CAVITY RINGDOWN ABSORPTION SPECTRUM OF THE T_{1}(n, π^{∗}) ← S_{0} TRANSITION OF ACROLEIN: ANALYSIS OF THE 0_{0}^{0} BAND ROTATIONAL CONTOUR

\underline{NIKOLAUS C. HLAVACEK}, MICHAEL O. MCANALLY, and STEPHEN DRUCKER, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702.

Acrolein (propenal, CH_{2}=CH—CH=O) is the simplest conjugated enal molecule and serves as a prototype for investigating the photochemical properties of larger enals and enones. Acrolein has a coplanar arrangement of heavy atoms in its ground electronic state. Much of the photochemistry is mediated by the T_{1}(π, π^{∗}) state, which has a CH_{2}–twisted equilibrium structure. In solution, the T_{1}(π, π^{∗}) state is typically accessed via intersystem crossing from an intially prepared planar S_{1}(n, π^{∗}) state. An intermediate in this photophysical transformation is the lowest ^{3}(n, π^{∗}) state, a planar species with adiabatic excitation energy below S_{1} and above T_{1}(π, π^{∗}). The present work focuses on this ^{3}(n, π^{∗}) intermediate state; it is designated T_{1}(n, π^{∗}) as the lowest-energy triplet state of acrolein having a planar equilibrium structure.

The T_{1}(n, π^{∗}) ← S_{0} band system, with origin near 412 nm, was first recorded in the 1970s at medium (0.5 cm^{−1}) resolution using a long-path absorption cell. Here we report the cavity ringdown spectrum of the 0_{0}^{0} band, recorded using a pulsed dye laser with 0.1 cm^{−1} spectral bandwidth. The spectrum was measured under both bulk-gas (roomtemperature) and jet-cooled conditions. The band contour in each spectrum was analyzed by using a computer program developed^{a} for simulating and fitting the rotational structure of singlet-triplet transitions. The assignment of several resolved sub-band heads in the room-temperature spectrum permitted approximate fitting of the inertial constants for the T_{1}(n, π^{∗}) state. The determined values (cm^{−1}) are A = 1.662, B = 0.1485, C = 0.1363. For the parameters A and (B + C)/2, estimated uncertainties of ±0.003 cm^{−1} and ±0.0004 cm^{−1}, respectively, correspond to a range of values that produce qualitatively satisfactory global agreement with the observed room-temperature contour. The fitted inertial constants were used to simulate the rotational contour of the 0_{0}^{0} band under jet-cooled conditions. Agreement with the observed jet-cooled spectrum was optimized by varying the homogeneous linewidth of the rovibronic transitions as well as the rotational temperature. The optimal FWHM was about 0.20 cm^{−1}, leading to an estimate of 25 ps for the lifetime of the T_{1}(n, π^{∗}) state of acrolein (v = 0) under isolated-molecule conditions.

aR. H. Judge et al., J. Chem. Phys. 103, 5343 (1995).